OpenStudy (anonymous):

Solve for s 6q + 2/9r = 6q +1/9s 2r + 12q, right?

OpenStudy (anonymous):

Let me have a look...

OpenStudy (anonymous):

Can't see how u got that, sorry.

OpenStudy (anonymous):

then wat's the answer?

OpenStudy (anonymous):

is this the question? \[\frac{6q+2}{9r} = \frac{6q+1}{9s}\]

OpenStudy (anonymous):

no its

OpenStudy (anonymous):

\[6q+\frac{2}{9r} = 6q +\frac{1}{9s}\] ?

OpenStudy (anonymous):

right answer is : 8s+9r=0

OpenStudy (anonymous):

yes! tht's it, joe!!!!

OpenStudy (anonymous):

oh the question is that which joe wrote there???

OpenStudy (anonymous):

alright, so first, because there is a 6q on both sides of the equation, we can subtract 6q from both sides and get rid of them: \[6q+\frac{2}{9r} = 6q+\frac{1}{9s} \Rightarrow \frac{2}{9r} = \frac{1}{9s}\] Does that make sense to you?

OpenStudy (anonymous):

yesh

OpenStudy (anonymous):

the final answer is 9r=18s ==> 9r-18s=0

OpenStudy (anonymous):

nice, now that we have gotten rid of those 6q's, we just have fractions left, and we can cross multiply:\[\frac{2}{9r} = \frac{1}{9s} \Rightarrow 2*9s = 1*9r \Rightarrow 18s = 9r\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The last thing we need to do in order to solve for s is to get rid of that 18 that is multiplying the s. So we divide both sides by 18: \[18s = 9r \Rightarrow s = \frac{9r}{18}\] then we reduce that fraction for the final answer: \[s = \frac{9r}{18} \Rightarrow s = \frac{r}{2}\]

OpenStudy (anonymous):

writing it like that or like: \[\frac{1}{2}r\] is acceptable. If there are any steps you are confused with or want me to explain further let me know :)